OpenStudy (anonymous):

Find the critical points for each function. Use the derivative test to determine whether the critical point is a local maximum, local minimum, or neither. y = x^4 - 8x^2

OpenStudy (campbell_st):

find the 1st and 2nd derivatives f'(x) = 4x^3 - 16x f"(x) = 12x^2 - 16 let f'(x) = 0 to find the stationary points 4x^3 - 16x = 0 4x(x^2 - 4) = 0 4x(x+2)(x-2)= 0 then startionary points at x = -2, 0, 2

OpenStudy (anonymous):

Right, thats where I got up to but I don't know where to go from there

OpenStudy (anonymous):

The answers are local minima: (-2, -16), (2,-16) local maximum: (0,0)

OpenStudy (campbell_st):

use the 2nd derivative test now f'(c) = 0 and f"(c) > 0 then y = f(x) has a minimum turning point at x = c f'(c) = 0 and f"(c) < 0 then y = f(x) has a maximum turning point at x = c f'(c) = 0 and f"c) = 0 then y = f(x) has a horizontal point of inflexion at x = c find f"(-2) = 48 - 16 >0 so a minimum f"(0) = - 16 < 0 so a maximum f"(2) = 48 - 16 > 0 so minimum

OpenStudy (campbell_st):

a point(s) of inflexion may exist at f"(x) = 0 so 12x^2 = 16 \[x = \pm 2/\sqrt{?}\]

OpenStudy (campbell_st):

?= 3